# Improving Multi-label Recognition using Class Co-Occurrence Probabilities

### v1

@@ Abstract
Multi-label Recognition (MLR) involves the identification of multiple objects within an image. To address the additional complexity of this problem, recent works have leveraged information from vision-language models (VLMs) trained on large text-images datasets for the task. These methods learn an independent classifier for each object (class), overlooking correlations in their occurrences. Such co-occurrences can be captured from the training data as conditional probabilities between a pair of classes. We propose a framework to extend the independent classifiers by incorporating the co-occurrence information for object pairs to improve the performance of independent classifiers. We use a Graph Convolutional Network (GCN) to enforce the conditional probabilities between classes, by refining the initial estimates derived from image and text sources obtained using VLMs. We validate our method on four MLR datasets, where our approach outperforms all state-of-the-art methods.
@@ 1 Introduction
Several approaches [41, 9] have taken the latter approach, namely recognition from image level features. Further, these approaches also reduce complexity by recognizing the presence of each object in an image independent of the others that may also be present. Training here amounts to learning independent classifiers for each object, which are then used to detect the corresponding objects, both using image level features. Apart from not segmenting the image and features into those for different objects, these methods also neglect the evidence for the presence of an object provided by the context of other objects. In practice, many objects are in sets, making their occurrences interdependent. Using independent classifiers neglects the mutual information present, which, if used, could enhance the performance of individual classifiers. This is particularly important in view of the already larger amount of data needed for MLR, and the relatively small sizes (vs SLR) of available annotated MLR datasets (because multiple classes present in an image require more annotation effort than required for SL images).
@@
We propose a two-stage framework that leverages the knowledge of VLMs but injects the co-occurrence information into the models. We obtain an initial estimate of the evidence logits for each class in a subimage in terms of the match between the embeddings of the subimage and those of the positive and negative text prompts associated with the class. Results from the subimages are aggregated to obtain logits for all classes across the image. These independently extracted subimage logits are refined to enforce prior knowledge of the joint probabilities of pairs of classes present in different subimages. Given that a class is present in a training image, the conditional probabilities of other classes being also present are measured from the frequencies of observing those classes in the image. Since the prior probability of a class is known from the VLM output, we can combine it with the conditional probabilities of other classes to obtain joint probabilities of class pairs. Enhancement of the outputs of the independent classifiers by using conditional probabilities is carried out through a graph convolutional network (GCN). GCN thus utilizes the conditional probabilities to improve upon the outputs of the independently obtained class estimates using the vision-language model. As commonly done, we compensate for the differences in the frequencies of different classes occurring in the training images by reweighting the estimated probabilities of different classes to remove the class bias in the training data.
@@
  • We propose a two-stage framework to adapt VLMs for MLR with limited annotated data, by enhancing the VLM based independent class estimates obtained in the first stage, with conditional probability priors extracted from the dataset, using a GCN.
@@ 2.1 Multi-Label Recognition
Multi-label recognition is an important, well-studied problem in computer vision, with a wide range of approaches being proposed to tackle it [32]. An important line of work has focused on learning disjoint binary classifiers for identifying each class of objects in an image [31, 37, 11, 9]. These approaches require large labeled datasets for training. They make no use of information that can be derived from modeling the co-occurrence of different classes, which is especially important when only a small amount of annotated data is available for training.
Other works have proposed to use recurrent neural networks (RNNs) to model label dependencies in an image [42, 30, 48]. Specifically, they cast MLR into a sequence prediction problem, using beam search to find a sequence of objects having the highest likelihood of being present in the image. Like our method, these methods also model label dependencies, but do so implicitly via the hidden state of RNNs. This requires vast amounts of labeled data to learn them. Recent approaches for MLR have proposed the use of VLMs to improve MLR performance when only a limited amount of labeled data is available.
@@ 2.2 Vision-Language Models for MLR
VLMs learn representations that are transferable to a wide range of downstream tasks such as classification [21, 43, 47, 51], retrieval [3, 26], and segmentation[45, 49, 50] by aligning hundreds of millions of image-text(prompts) pairs. Such approaches commonly focus on learning prompts suitable for these downstream tasks [53]. [41] adapts VLMs for MLR, proposing to learn a pair of prompts associated with the presence/ absence of each class. Text embeddings extracted from the learned prompts are used to gather local evidence from the image features extracted by the VLM, which is then aggregated and combined. On the other hand, SCPNet [17] is another VLM based approach for MLR which uses the similarity between the class prompt embedding from VLMs to train a classifier for each class. In these methods, classifiers for each class are independent of those for other classes, not making use of any contextual information during inference on an image.
@@ 3 Method
Figure 1: Method Overview: Given an image with multiple objects, we extract image features and text features from the subimages using a vision-language model (CLIP). An image-text feature aggregation module (Sec. 3.1) combines these features to identify all classes present in the image as a union of the classes present in the subimages, giving an initial set of image level class logits. These logits are passed to a GCN, that uses conditional probabilities between classes to refine these initial predictions (Sec. 3.2). We train this framework while reweighting the loss generated by classes to address any class imbalance in the training data using a Reweighted Asymmetric Loss (RASL), a weighted version of the familiar ASL.
@@ 4.4 Results
We compare our approach primarily with DualCoOp [41] and SCPNet [17] which use VLMs like our method, on all four standard benchmarks discussed earlier. As seen in Table 1, our method outperforms DualCoOp by 0.4% and SCPNet by 7.0% mAP on the VOC-2007. On COCO-small, our method outperforms DualCoOp by 2.4% and SCPNet by 3.3% mAP.
On the FoodSeg103 dataset, our approach significantly improves upon DualCoOp by 3.9% and SCPNet by 4.1% mAP. In the UNIMIB dataset, our method achieves substantial performance gains of 14.1% over DualCoOp and 12.2% mAP over SCPNet.
@@ 4.6 Performance on Classes that are Difficult to Recognize
Table 2 compares the performance of our method on these classes with the previous state-of-the-art DualCoOp[41]. We see that our method significantly improves the performance of DualCoOp[41], which relies solely on VLMs without modeling any conditional probabilities. Specifically, we observe a growth of 22.3% in CP, 33.9% in CR and 34.8% for CF1 on UNIMIB2016. On FoodSeg103, our method significantly improves CP, CR and CF1, by 15.6%, 7.2% and 11.89%, respectively. This underscores the importance of the information obtained by modeling joint class probabilities in recognizing classes of objects that are difficult to recognize from image features alone.
@@
DualCoOp Ours w/o reweigh Ours DualCoOp Ours w/o reweigh Ours
CP 25.4 41.9 57.6 13.7 14.8 28.7
CR 26.2 57.5 60.0 19.7 22.5 26.9
CF1 24.3 44.9 59.1 16.5 18.7 28.4
Table 2: A comparison of the average performance of our approach with the previous state-of-the-art VLM-based method DualCoOp[41] on classes that are difficult to recognize using only visual features (having 10 lowest CF1 values on the FoodSeg103[44] and UNIMIB[10]). Our approach significantly improves MLR performance on such classes due to its use of information derived from class conditional probabilities.
@@ 7 Acknowledgement
The support of the Office of Naval Research under grant N00014-20-1-2444, of USDA National Institute of Food and Agriculture under grant 2020-67021-32799/1024178 and Vizzhy.com are gratefully acknowledged.

### v2

@@ Abstract
Multi-label Recognition (MLR) involves the identification of multiple objects within an image. To address the additional complexity of this problem, recent works have leveraged information from vision-language models (VLMs) trained on large text-image datasets for the task. These methods learn an independent classifier for each object (class), overlooking correlations in their occurrences. Such co-occurrences can be captured from the training data as conditional probabilities between a pair of classes. We propose a framework to extend the independent classifiers by incorporating the co-occurrence information for object pairs to improve the performance of independent classifiers. We use a Graph Convolutional Network (GCN) to enforce the conditional probabilities between classes, by refining the initial estimates derived from image and text sources obtained using VLMs. We validate our method on four MLR datasets, where our approach outperforms all state-of-the-art methods.
@@ 1 Introduction
Several approaches [42] have taken the latter approach, namely recognition from image level features. Further, these approaches also reduce complexity by recognizing the presence of each object in an image independent of the others that may also be present. Training here amounts to learning independent classifiers for each object, which are then used to detect the corresponding objects, both using image level features. Apart from not segmenting the image and features into those for different objects, these methods also neglect the evidence for the presence of an object provided by the context of other objects. In practice, many objects are in sets, making their occurrences interdependent. Using independent classifiers neglects the mutual information present, which, if used, could enhance the performance of individual classifiers. This is particularly important in view of the already larger amount of data needed for MLR, and the relatively small sizes (vs SLR) of available annotated MLR datasets (because multiple classes present in an image require more annotation effort than required for SL images).
@@
We propose a two-stage framework that leverages the knowledge of VLMs but injects the co-occurrence information into the models. We obtain an initial estimate of the evidence logits for each class in a subimage in terms of the match between the embeddings of the subimage and those of the positive and negative text prompts associated with the class. Results from the subimages are aggregated to obtain logits for all classes across the image. These independently extracted subimage logits are refined to enforce prior knowledge of the joint probabilities of pairs of classes present in different subimages. Given that a class is present in a training image, the conditional probabilities of other classes being also present are measured from the frequencies of observing those classes in the image. Since the prior probability of a class is known from the VLM output, we can combine it with the conditional probabilities of other classes to obtain joint probabilities of class pairs. Enhancement of the outputs of the independent classifiers by using conditional probabilities is carried out through a graph convolutional network (GCN). GCN thus utilizes the conditional probabilities to improve upon the outputs of the independently obtained class estimates using the vision-language model. As commonly done, we compensate for the differences in the frequencies of different classes occurring in the training images by reweighing the estimated probabilities of different classes to remove the class bias in the training data.
@@
  • We propose a two-stage framework to adapt VLMs for MLR with limited annotated data, by enhancing the VLM-based independent class estimates obtained in the first stage, with conditional probability priors extracted from the dataset, using a GCN.
@@ 2.1 Multi-Label Recognition
Multi-label recognition is an important, well-studied problem in computer vision, with a wide range of approaches being proposed to tackle it [33]. An important line of work has focused on learning binary classifiers for identifying each class of objects in an image [13, 38]. These approaches require large labeled datasets for training. They make no use of information that can be derived from modeling the co-occurrence of different classes, which is especially important when only a small amount of annotated data is available for training.
Other works attempt to model label dependencies in an image. [11] models label dependencies implicitly by using a deep embedding network shared by classifiers for all different classes of objects. Other works have proposed to use recurrent neural networks (RNNs) to model label dependencies in an image [32, 43, 49]. Specifically, they cast MLR into a sequence prediction problem, using beam search to find a sequence of objects having the highest likelihood of being present in the image. Like our method, these methods also model label dependencies, but do so via the hidden state of RNNs. However, both these approaches: a deep embedding network or an RNN require large amounts of labeled data for training.
[11] also proposed to guide the training of the deep embedding network using a GCN that explicitly models the label correlation matrix. However, during inference, only the correlation implicitly learned by the deep embedding network is used for MLR. In contrast, we explicitly enforce label correlations using a GCN during both training and inference. We also note that our approach does not rely on implicitly learning label dependencies using the embedding network, as it remains frozen to preserve VLM priors.
@@ 2.2 Vision-Language Models for MLR
VLMs learn representations that are transferable to a wide range of downstream tasks such as recognition [5, 23, 44, 48], retrieval [3, 4, 28], and segmentation[46, 50, 51] by aligning hundreds of millions of image-text(prompts) pairs. Such approaches commonly focus on learning prompts suitable for these downstream tasks [53]. [42] adapts VLMs for MLR, proposing to learn only a pair of prompts associated with the presence/ absence of each class while keeping the VLM itself frozen. Text embeddings extracted from the learned prompts are used to gather local evidence from the image features extracted by the VLM, which is then aggregated and combined. SCPNet [19] is another VLM based approach for MLR, differing from [42] in its use of a GCN to help learn text embeddings for the presence/absence prompts. The GCN models priors derived from class name similarities in CLIP’s embedding space, enforcing these priors during prompt learning. [19] also augments training with a self supervised contrastive loss. Both these methods learn independent prompt-based classifiers (as VLM is frozen), not making any use of co-occurrence information during inference. Our method instead learns inter-dependent classifiers by using a GCN to enhance predictions made by the VLM. The inter-dependent classifiers model and make use of conditional probabilities during both training and inference and use co-occurrence information from the actual training dataset.
@@ 3 Method
Figure 1: Method Overview: Given an image with multiple objects, we extract image features and text features from the subimages using a vision-language model (CLIP). An image-text feature aggregation module (Sec. 3.1) combines these features to identify all classes present in the image as a union of the classes present in the subimages, giving an initial set of image level class logits. These logits are passed to a GCN, that uses conditional probabilities between classes to refine these initial predictions (Sec. 3.2). We train this framework while reweighting the loss generated by classes to address any class imbalance in the training data using a Reweighted Asymmetric Loss (RASL), a weighted version of ASL[41].
@@ 4.4 Results
We primarily compare our approach with DualCoOp [42] and SCPNet [19], as they are the only other MLR baselines that use VLMs, making them SOTA methods in limited data settings across all four standard benchmarks discussed earlier. As seen in Table 1, our method outperforms DualCoOp by 0.4% and SCPNet by 7.0% mAP on the VOC-2007. On COCO-small, our method outperforms DualCoOp by 2.4% and SCPNet by 3.3% mAP. On the FoodSeg103 dataset, our approach significantly improves upon DualCoOp by 3.9% and SCPNet by 4.1% mAP. In the UNIMIB dataset, our method achieves substantial performance gains of 14.1% over DualCoOp and 12.2% mAP over SCPNet.
@@ 4.6 Performance on Classes that are Difficult to Recognize
Table 2 compares the performance of our method on these classes with the previous SOTA DualCoOp[42] and SCPNet[19]. We see that our method significantly improves the performance of these methods, which relies solely on VLMs without modeling any conditional probabilities. Specifically, for DualCoOp, we observe a growth of 22.3% in CP, 33.9% in CR and 34.8% for CF1 on UNIMIB2016, and 15.6% in CP, 7.2% in CR and 11.89% in CF on FoodSeg103. For SCPNet, we see gains of 27.1% CP, 25.2% CR, and 26.5% CF1 on UNIMIB2016, and 15.8% CP, 5.8% CR, and 12.4% CF1 on FoodSeg103.
This underscores the importance of the information obtained by modeling joint class probabilities in recognizing classes of objects that are difficult to recognize from image features alone.
@@
Methods CP CR CF1 CP CR CF1
DualCoOp 25.4 26.2 24.3 13.7 19.7 16.5
SCPNet 30.5 34.8 32.5 12.9 21.1 16.0
Ours w/o reweigh 41.9 57.5 44.9 14.8 22.5 18.7
Ours 57.6 60.0 59.1 28.7 26.9 28.4
Table 2: A comparison of the average performance of our approach with the previous state-of-the-art VLM-based method DualCoOp[42] and SCPNet[19] on classes that are difficult to recognize using only visual features (having 10 lowest CF1 values on the FoodSeg103[45] and UNIMIB[12]). Our approach significantly improves MLR performance on such classes due to its use of information derived from class conditional probabilities.
@@ 7 Acknowledgement
We thank Kamila Abdiyeva for her insightful feedback on the manuscript. The support of the Office of Naval Research under grant N00014-20-1-2444, of USDA National Institute of Food and Agriculture under grant 2020-67021-32799/1024178 and Vizzhy.com are gratefully acknowledged.